OpenStudy (anonymous):

What is the sum of the angle measures in a polygon that has 20 sides? A.2,000° B.3,240° C.3,600° D.4000°

OpenStudy (anonymous):

360/20=18 180-18=162 162*20=3.240 is C medal help ?

OpenStudy (anonymous):

What is the measure of each angle in a regular polygon with 18 sides? a.156 b.160 c.180 d.320